OpenStudy (anonymous):

What is the x-coordinate of the vertex of y = |4x + 12| + 3 ? –3 0 3 4

OpenStudy (anonymous):

@Mitch2014 @mathslover @nitz

mathslover (mathslover):

It will be that value for which the absolute term is zero.

mathslover (mathslover):

that is, 4x + 12 = 0

mathslover (mathslover):

Solve for x now.

OpenStudy (anonymous):

if its 4x+12=0 @mathslover ima do 4x+12=-3 Am i right

OpenStudy (anonymous):

@RyanL.

OpenStudy (anonymous):

@mathslover Vertex is where the sign of "inside expression" inside the absolute brackets changes sign, that is passes through zero. Itmeans the vertex is where 4x+12=0

mathslover (mathslover):

Yeah. , you get : 4x + 12 = 0

mathslover (mathslover):

I said the same...

OpenStudy (anonymous):

:O sorry @mathslover thnx

mathslover (mathslover):

You're welcome barbie :)